OpenStudy (adamk):

A 0.20-kg particle moves along the x axis under the in°uence of a stationary object. The potential energy is given by U(x) = (8.0 J/m^2)x^2 + (2.0 J/m^4)x^4 ; where x is in coordinate of the particle. If the particle has a speed of 5:0 m=s when it is at x = 1:0 m, its speed when it is at the origin is: A. 0 B. 2:5 m=s C. 5:7 m=s D. 7:9 m=s E. 11 m=s ans: E Not sure how to do this.

OpenStudy (raffle_snaffle):

You sure this isn't kinetic energy? Potential energy is in terms of other variables, not velocity.

OpenStudy (adamk):

@raffle_snaffle I figured it out. You basically just use KE + PE = constant to solve.

OpenStudy (raffle_snaffle):

Cool

OpenStudy (irishboy123):

Generally for a conservative energy field function \(U\), so you are assuming that it is a conservative field in this case, there is always a vector \(\vec F\) such that, \(\vec F = - \nabla U\), and for motion \(\vec F = m \vec {\ddot x }\) in a 1-D system that simplifies to: \(F= m \ddot x = - \dfrac{dU}{dx}\) Write it as \(m \dot x \dfrac{d \dot x}{dx} = - \dfrac{dU}{dx}\) Integrate: \(\frac{1}{2} m \dot x^2 = - U + C\) Apply the IV: \(\frac{1}{2} (0.2) 5^2 = - 10 + C \implies C = 12.5 \) \(\frac{1}{2} m \dot x^2 = - U(x) + 12.5\) \( \dot x = \sqrt{ \dfrac{ 25 - 2U(x)}{ m} }\) \( \dot x(0) = \sqrt{ \dfrac{ 25 - 0}{ 0.2} } = 11.18\) m/s

OpenStudy (irishboy123):

And that was completely OTT as i realised when i pushed the send button :( once you make the assumption it's conservative, you write and solve \(\frac{1}{2} m \dot x^2 + U = C\) where C is the constant of the motion, in this case total energy you then own it, as you did